Imagine that 2 genes code for stripes in canines. 2 wolves have pups that have the phenotypic ratio as follows : 12/16 stripes, 3/16 semi stripes and 1/16 no stripes. What is the genotype of the no stripes progeny?
a. bb_cc
b. BbC_
c. B_cc
d.B_C_ and B_cc
e. B_C_
f. None of the above
Answer #1

Answer:

a). bbcc

Explanation:

B_C_ + B_cc = 12/16 (stripes)

bbC_ = 3/16 (semi stripes)

bbcc = 1/16 (no stripes)